OpenStudy (anonymous):

Evaluate: 4a+7b+3a-2b for a=5 b=-3

OpenStudy (anonymous):

Just plug in 5 for every "a" and -3 for every "b"

OpenStudy (anonymous):

4(5) + 7(-3) + 3(5) - 2(-3)

OpenStudy (anonymous):

is it 20 or 8?

OpenStudy (anonymous):

Break it up into four parts: 4(5) = 20 7(-3) = -21 3(5) = 15 2(-3) = -6 So, your equation is now: 20 + (-21) + 15 - (-6)

OpenStudy (anonymous):

so its 20 right {;

OpenStudy (anonymous):

20 + (-21) = -1 -1 + 15 = 14 14 - (-6) = 14 + 6 = 20 Looks about right.